# Simulated Annealing for Quadratic and Higher-Order Unconstrained Integer Optimization

### v1

@@ Abstract
Simulated annealing (SA) is a key algorithm for solving combinatorial optimization problems, which model numerous real-world systems. While SA is commonly used to solve quadratic unconstrained binary optimization (QUBO) problems, many practical problems are more naturally formulated using integer variables. We therefore study quadratic and higher-order unconstrained integer optimization (QUIO and HUIO) problems, which generalize QUBO by allowing integer-valued variables and higher-order interactions. Conventional approaches often convert these problems into QUBO formulations through binary encoding and the reduction of higher-order terms. Such conversions, however, greatly increase the number of variables and interactions, resulting in long computation times and, for large-scale problems, even making the conversion itself a dominant computational bottleneck. To overcome this limitation, we propose an efficient framework that directly applies SA to QUIO and HUIO problems without converting them into QUBO. Within this framework, we introduce an optimal-transition Metropolis method, designed to improve efficiency when the variable bounds are wide, and evaluate its performance alongside the conventional Metropolis and heat bath methods. Numerical experiments demonstrate that the proposed direct approach achieves higher efficiency and solution quality than the conventional QUBO-based formulation and reveal the practical advantages of the optimal-transition Metropolis method. The algorithm developed in this study is available as part of the open-source library OpenJij, which provides a Python interface with a C++ backend.
@@ 1 Introduction
Combinatorial optimization problems play a central role in a wide range of real-world applications, including finance, logistics, drug discovery, and scheduling. Many of these problems are NP-complete or NP-hard, and obtaining exact solutions in polynomial time is generally intractable.[fphy.2014.00005] Therefore, efficient approximation algorithms, often referred to as metaheuristics, are indispensable. Among them, simulated annealing (SA)[SA, SA2], a highly versatile metaheuristic inspired by the analogy with the physical annealing process, is one of the most powerful and widely used approaches due to its ability to escape from local optima and explore the global optimum.
@@
When applying SA to combinatorial optimization problems, it is common to reformulate them as quadratic unconstrained binary optimization (QUBO) problems.[fphy.2014.00005, Kochenberger2014, Glover2022] A QUBO problem is defined as minimizing a quadratic polynomial whose variables take binary values $\{0,1\}$. However, in many real-world cases, modeling the problem using integer variables is more natural than using binary ones. In such cases, it is natural to formulate the problem as a quadratic unconstrained integer optimization (QUIO) problem. Furthermore, these problems often contain higher-order polynomial terms or constraints, which can be represented by extending QUIO to the higher-order unconstrained integer optimization (HUIO) formulation.
@@
In previous studies, QUIO and HUIO problems have often been solved by converting them into QUBO formulations. However, this conversion inherently introduces two major computational bottlenecks. First, binary encoding is required to transform each integer variable into multiple binary variables.[Karimi2019, 9435359] Second, when the objective function contains higher-order terms beyond quadratic, higher-order term reduction must be performed to express them in quadratic form.[BOROS2002155, Anthony2017] These conversion procedures dramatically increase both the number of variables and the number of interactions, thereby making the resulting QUBO model denser and its search space more complex. For large-scale problems, the conversion process itself can become a dominant computational overhead, and in some cases it does not finish within a practical amount of time, which makes it a serious limitation in practice. Therefore, it is desirable to solve such problems without converting them into QUBO. Although SA can in principle be applied directly to QUIO and HUIO formulations, actual applications of this approach remain scarce. Existing studies address only limited classes of QUIO problems[CARDOSO19971349, Abramson1999, integer_sa_1], and research on the general HUIO case appears to be extremely scarce.
@@
In this study, we propose an efficient framework that directly applies SA to general QUIO and HUIO formulations without converting them into QUBO, and we describe in detail how SA can be implemented efficiently for such problems. In typical SA implementations for QUBO problems, a single-spin flip update, where one binary variable is flipped at a time, is commonly used. Similarly, in our framework, a single-variable update is employed, where only one integer variable is changed at a time. Under this scheme, the energy difference required for variable updates, which is often a computational bottleneck in SA, can be efficiently computed even for QUIO and HUIO formulations. As for transition probabilities, the Metropolis method[METROPOLIS, METROPOLIS_HASTING] and the heat bath method[10.1071/PH650119, 4767596] are widely used; however, when the bounds of integer variables are large, both approaches encounter difficulties. To overcome this issue, we introduce the optimal-transition Metropolis method, which enables efficient updates under wide variable bounds. Using these transition schemes, we apply SA directly to QUIO and HUIO and evaluate their performance, including comparisons with the conventional QUBO-based approach.
@@
This paper is organized as follows. Section 2 presents the formulation of QUIO and HUIO. Section 3 describes the proposed SA algorithm in detail, with particular emphasis on the computation of energy differences and transition probabilities. Section 4 reports the numerical results, and Section 5 concludes the paper with a discussion of future perspectives.
@@ 2 Model
In this section, we introduce the polynomial unconstrained integer optimization, QUIO and HUIO. We begin with the quadratic unconstrained integer optimization. QUIO is defined as the problem of minimizing a quadratic function composed of integer variables:
@@
In the case of binary variables, HUIO reduces to the higher-order unconstrained binary optimization (HUBO) problems.[IJMHEUR.2011.041196, IJMHEUR.2011.044356] HUBO formulations also arise naturally in real-world optimization models, and there are studies in which SA has been applied directly to such cases.[PhysRevA.109.032416, Wang2025, ikeuchi2025] HUIO can be viewed as a natural extension of QUIO or HUBO, and higher-order terms of this form naturally arise in practical optimization problems, for example when handling quadratic constraints involving integer variables using penalty or augmented Lagrangian methods.[Bertsekas1982]
In the context of mathematical optimization, $E_{p}(\bm{z})$ is referred to as the objective function, and its minimum value is called the optimal value, while any configuration $\bm{z}$ that gives this value is called an optimal solution. In general, multiple optimal solutions may exist that yield the same optimal value. From the viewpoint of statistical physics, $E_{p}(\bm{z})$ corresponds to the Hamiltonian of a classical spin system, where the variables $z_{i}$ can be regarded as generalized spin variables not restricted to the usual $S=1/2$ case. In this interpretation, QUIO and HUIO represent the problem of finding the ground state of the spin system and its corresponding energy, and the existence of multiple optimal solutions corresponds to degeneracy of the ground state. In this paper, we use the terminology from optimization and physics appropriately depending on the context.
@@ 3.1 Algorithm
    Evaluate the transition probability $P(\Delta E_{p,k},T)$ and accept the change $z_{k}+\Delta z_{k}$ according to this probability.
@@ 3.2 Energy difference
To perform SA efficiently, the energy difference when a variable moves must be calculated efficiently. A common method for QUBO problems is to precompute the energy difference for the flip of each variable and update these values when a flip is accepted. In this approach, the energy difference can be obtained in $O(1)$ time, which is particularly advantageous when variable updates are infrequent. In fact, during the low-temperature stage of SA, the system is already close to its ground state and variable updates rarely occur. As a result, this method improves the computational efficiency.
However, this method is only applicable to binary variables, as they can transition to only one alternative state. For QUIO and HUIO, an integer variable may transition to one of multiple possible states, and thus this strategy cannot be applied directly. Instead of precomputing the energy differences themselves, we store the corresponding coefficients.
@@ 3.3.1 Metropolis method
The advantage of this update method is that the determination of the variable change does not depend on the number of possible changes of the variable, $u_{k}-l_{k}+1$. In QUIO and HUIO problems arising in mathematical optimization, variables may have very wide bounds; nevertheless, the transition probability can still be evaluated in $O(1)$ time. On the other hand, since the variable difference is chosen at random, the probability of selecting the optimal update for a single variable $z_{k}$ scales as $O(1/(u_{k}-l_{k}))$ when the bounds are very wide, i.e., $u_{k}-l_{k}\gg 1$.
@@ 3.3.2 Heat bath method
Next, we describe the heat bath method, which is another widely used method. In this method, the probability of selecting the update with energy difference $\Delta E^{(i)}_{k}$ for $z_{k}\to z_{k}+\Delta z^{(i)}_{k}$ is given by
$P(\Delta E^{(i)}_{k},T)=\frac{\displaystyle\exp\left(-\frac{\Delta E^{(i)}_{k}}{T}\right)}{\displaystyle\sum^{u_{k}-l_{k}}_{i=0}\exp\left(-\frac{\Delta E^{(i)}_{k}}{T}\right)}.$ (16)
Unlike the Metropolis method, the variable difference is not chosen at random. Instead, the probability is determined directly by the corresponding energy differences $\Delta E^{(i)}_{k}$. For this reason, the heat bath method is often considered more efficient than the Metropolis method. However, it is important to note that in the heat bath method one must compute the transition probability for all possible changes $\Delta z^{(i)}_{k}$ in general. As a result, the computational cost scales as $O(u_{k}-l_{k})$, which can become prohibitive when the bounds of a variable are large, significantly increasing the overall runtime of SA.
@@ 3.3.3 Optimal-transition Metropolis method
Note that, if this transition were always chosen, the system would be trapped in a local minimum. To avoid this, we introduce a parameter $r=i/N_{\text{sweep}}$ ($i=1,2,\ldots,N_{\text{sweep}}$), defined by the progress of the annealing schedule, which controls how often the optimal transition is attempted. Namely, the probability that a candidate change $\Delta z_{k}^{(i)}$ is accepted is given by
@@
Here, $\Delta E^{\min}_{k}=\min_{\Delta z_{k}}\,\Delta E_{p,k}$ is the minimum energy difference among all possible changes of $z_{k}$. We also note that similar attempts to improve proposal mechanisms have been made in SA for continuous variables.[Dekkers1991, Ali1997, ALI200287, Locatelli2002] In general, obtaining $\Delta z_{\min}$ requires minimizing a polynomial expression of the energy difference with respect to $\Delta z$, which can be computationally expensive. However, when the polynomial degree is small, the minimum can be found efficiently. In this study, we implemented the optimal-transition Metropolis method in cases where the energy difference is at most a fourth-order polynomial in $\Delta z$, which corresponds to HUIO problems in which each variable appears with degree at most four within each monomial.
It should be noted that, unlike the Metropolis and heat bath methods, the optimal-transition Metropolis method does not satisfy the balance condition for the Boltzmann distribution even when the temperature $T$ and the parameter $r>0$ are kept fixed. Therefore, it cannot be used for equilibrium sampling at a fixed temperature and can be employed only as a heuristic for searching low-energy states within SA.
@@ 3.4 Initial and final temperature
In this subsection, we explain how the initial and final temperatures in SA are determined.[10.1063/1.34823, Aarts1985StatisticalC] The initial temperature must be sufficiently high for the system. If it is set too low, the system easily becomes trapped in local minima, and the quality of the final solution is degraded. Conversely, if it is set too high, the search process performed at such temperatures becomes essentially wasteful. In general, using a temperature that is too high is less problematic than using one that is too low. However, when the total annealing time is fixed, setting the temperature too high effectively shortens the low-temperature phase of the annealing process, which again reduces the quality of the final solution. Therefore, the initial temperature should be chosen to be sufficiently high but not excessive.
The final temperature must likewise be sufficiently low for the system. If it remains too high, the energy cannot decrease enough, and the quality of the final solution is also degraded. If it is too low, on the other hand, state updates rarely occur and computation time is wasted without further improvement, although setting it too low is still preferable to too high.
An important point is that the appropriate initial and final temperatures cannot be determined solely from the properties of the model. They also depend on other factors such as the total annealing time, the update scheme of the states, and the specific temperature schedule. However, it is quite difficult to determine these temperatures in a way that also accounts for these factors. In this study, therefore, we determine both temperatures solely based on the model information. To ensure robustness against variations in the annealing time or cooling schedule, the initial temperature is slightly overestimated on the high-temperature side, and the final temperature is slightly overestimated on the low-temperature side.
@@
$T_{\text{init}}=-\frac{\Delta E_{\text{max}}}{\log a}$ (26)
@@
$\tilde{a}^{(m)}_{k}(\bm{z})$ $=\|J_{\underbrace{k,\dots,k}_{m\text{ times}}}\|+\sum_{\begin{subarray}{c}i_{1}\\ i_{1}\neq k\end{subarray}}\|J_{i_{1},\underbrace{k,\dots,k}_{m\text{ times}}}\|z^{\text{max}}_{i_{i}}$
$+\sum_{\begin{subarray}{c}i_{1},i_{2}\\ i_{1}\neq k,i_{2}\neq k\end{subarray}}\|J_{i_{1},i_{2},\underbrace{k,\dots,k}_{m\text{ times}}}\|z^{\text{max}}_{i_{i}}z^{\text{max}}_{i_{2}}+\cdots,$ (28)
@@
$T_{\text{final}}=-\frac{\Delta E_{\min}}{\log b}.$ (30)
@@
These definitions of the initial and final temperatures provide a practical way to set appropriate temperature ranges for SA based only on the properties of the objective function.

### v2

@@ Abstract
Simulated annealing (SA) is a key algorithm for solving combinatorial optimization problems, which model numerous real-world systems. While SA is commonly used to solve quadratic unconstrained binary optimization (QUBO) problems, many practical problems are more naturally formulated using integer variables. We therefore study quadratic and higher-order unconstrained integer optimization (QUIO and HUIO) problems, which generalize QUBO by allowing integer-valued variables and higher-order interactions. Conventional approaches often convert these problems into QUBO formulations through binary encoding and the reduction of higher-order terms. Such conversions, however, greatly increase the number of variables and interactions, resulting in long computation times and, for large-scale problems, even making the conversion itself a dominant computational bottleneck. To overcome this limitation, we propose an efficient framework that directly applies SA to QUIO and HUIO problems without converting them into QUBO. Within this framework, we introduce an optimal-transition Metropolis method, designed to improve efficiency when the variable range is wide, and evaluate its performance alongside the conventional Metropolis and heat bath methods. Numerical experiments demonstrate that the proposed direct approach achieves higher efficiency and better solution quality than the conventional QUBO-based formulation and reveal the practical advantages of the optimal-transition Metropolis method. The algorithm developed in this study is available as part of the open-source library OpenJij, which provides a Python front end with a C++ backend.
@@ 1 Introduction
Combinatorial optimization problems play a central role in a wide range of real-world applications, including finance, logistics, drug discovery, and scheduling. Many of these problems are NP-complete or NP-hard, and obtaining exact solutions in polynomial time is generally intractable.[fphy.2014.00005] Therefore, efficient approximation algorithms, often referred to as metaheuristics, are indispensable. Among them, simulated annealing (SA)[SA, SA2], a highly versatile metaheuristic inspired by the analogy with the physical annealing process, is one of the most powerful and widely used approaches due to its ability to escape from local optima and search for the global optimum.
@@
When applying SA to combinatorial optimization problems, it is common to reformulate them as quadratic unconstrained binary optimization (QUBO) problems.[fphy.2014.00005, Kochenberger2014, Glover2022] A QUBO problem is defined as minimizing a quadratic polynomial whose variables take binary values $\{0,1\}$. However, in many real-world cases, modeling the problem with integer variables is more natural than with binary variables. In such cases, it is natural to formulate the problem as a quadratic unconstrained integer optimization (QUIO) problem. Furthermore, these problems often contain higher-order polynomial terms or constraints, which can be represented by extending QUIO to the higher-order unconstrained integer optimization (HUIO) formulation.
@@
In previous studies, QUIO and HUIO problems have often been solved by converting them into QUBO formulations. However, this conversion inherently introduces two major computational bottlenecks. First, binary encoding is required to transform each integer variable into multiple binary variables.[Karimi2019, 9435359] Second, when the objective function contains higher-order terms beyond quadratic, higher-order term reduction must be carried out to express them in quadratic form.[BOROS2002155, Anthony2017] These conversion procedures dramatically increase both the number of variables and the number of interactions, thereby making the resulting QUBO model much denser and its search space more complex. For large-scale problems, the conversion process itself can become a dominant computational overhead, and in some cases it does not finish within a practical amount of time, which makes it a serious limitation in practice. Therefore, it is desirable to solve such problems without converting them into QUBO. Although SA can in principle be applied directly to QUIO and HUIO formulations, actual applications of this approach remain scarce. Existing studies address only limited classes of QUIO problems[CARDOSO19971349, Abramson1999, integer_sa_1], and research on the general HUIO case appears to be extremely scarce.[TIAN1995629]
@@
In this study, we propose an efficient framework that directly applies SA to general QUIO and HUIO formulations without converting them into QUBO, and we describe in detail how SA can be implemented efficiently for such problems. In typical SA implementations for QUBO problems, a single-spin flip update, where one binary variable is flipped at a time, is commonly used. Similarly, in our framework, a single-variable update is employed, where only one integer variable is changed at a time. Under this scheme, the energy difference required for variable updates, which is often a computational bottleneck in SA, can be efficiently computed even for QUIO and HUIO formulations. As for transition probabilities, the Metropolis method[METROPOLIS, METROPOLIS_HASTING] and the heat bath method[10.1071/PH650119, 4767596] are widely used; however, when the range of integer variables are large, both approaches encounter difficulties. To overcome this issue, we introduce the optimal-transition Metropolis method, which enables efficient updates under wide variable range. Using these transition schemes, we apply SA directly to QUIO and HUIO and evaluate their performance, including comparisons with the conventional QUBO-based approach.
@@
This paper is organized as follows. Section 2 presents the formulations of QUIO and HUIO. Section 3 describes the proposed SA algorithm in detail, with particular emphasis on the computation of energy differences and transition probabilities. Section 4 reports the numerical results, and Section 5 concludes the paper with a discussion of future perspectives.
@@ 2 Model
In this section, we introduce the polynomial unconstrained integer optimization, namely QUIO and HUIO. We begin with the quadratic unconstrained integer optimization. QUIO is defined as the problem of minimizing a quadratic function composed of integer variables:
@@
In the case of binary variables, HUIO reduces to the higher-order unconstrained binary optimization (HUBO).[IJMHEUR.2011.041196, IJMHEUR.2011.044356] HUBO formulations also arise naturally in real-world optimization models, and there are studies in which SA has been applied directly to such cases.[PhysRevA.109.032416, Wang2025, ikeuchi2025] HUIO can be viewed as a natural extension of QUIO or HUBO, and higher-order terms of this form naturally arise in practical optimization problems, for example when handling quadratic constraints involving integer variables using penalty or augmented Lagrangian methods.[Bertsekas1982]
In the context of mathematical optimization, $E_{p}(\bm{z})$ is referred to as the objective function, and its minimum value is called the optimal value, while any configuration $\bm{z}$ that gives this value is called an optimal solution. In general, multiple optimal solutions may exist that yield the same optimal value. From the viewpoint of statistical physics, $E_{p}(\bm{z})$ corresponds to the Hamiltonian of a classical spin system, where the variables $z_{i}$ can be regarded as generalized classical spin variables that take arbitrary integer values. In this interpretation, QUIO and HUIO represent the problem of finding the ground state of the spin system and its corresponding energy, and the existence of multiple optimal solutions corresponds to degeneracy of the ground state. In this paper, we use the terminology from optimization and physics appropriately depending on the context.
@@ 3.1 Algorithm
    Evaluate the transition probability $P(\Delta E_{p,k},T)$ and accept the change $z_{k}\to z_{k}+\Delta z_{k}$ according to this probability.
@@ 3.2 Energy difference
To perform SA efficiently, the energy difference when a variable changes must be calculated efficiently. A common method for QUBO problems is to precompute the energy difference for the flip of each variable and update these values when a flip is accepted. In this approach, the energy difference can be obtained in $O(1)$ time, which is particularly advantageous when variable updates are infrequent. In fact, during the low-temperature stage of SA, the system is already close to its ground state and variable updates rarely occur. As a result, this method improves the computational efficiency.
However, this method is only applicable to binary variables, because they can transition to only one alternative state. For QUIO and HUIO, an integer variable may transition to one of multiple possible states, and thus this strategy cannot be applied directly. Instead of precomputing the energy differences themselves, we store the corresponding coefficients.
@@ 3.3.1 Metropolis method
The advantage of this update method is that the determination of the variable change does not depend on the number of possible changes of the variable, $u_{k}-l_{k}+1$. In QUIO and HUIO problems arising in mathematical optimization, variables may have very wide range; nevertheless, the transition probability can still be evaluated in $O(1)$ time. On the other hand, since the variable difference is chosen at random, the probability of selecting the optimal update for a single variable $z_{k}$ is $O(1/(u_{k}-l_{k}))$, and becomes very small when the variable ranges are wide.
@@ 3.3.2 Heat bath method
Next, we describe the heat bath method, which is another widely used method. In this method, the probability of taking the update with energy difference $\Delta E^{(i)}_{k}$ for $z_{k}\to z_{k}+\Delta z^{(i)}_{k}$ is given by
$P(\Delta E^{(i)}_{k},T)=\frac{\displaystyle\exp\left(-\frac{\Delta E^{(i)}_{k}}{T}\right)}{\displaystyle\sum^{u_{k}-l_{k}}_{j=0}\exp\left(-\frac{\Delta E^{(j)}_{k}}{T}\right)}.$ (16)
Unlike the Metropolis method, the variable difference is not chosen at random. Instead, the probability is determined directly by the corresponding energy differences $\Delta E^{(i)}_{k}$. For this reason, the heat bath method is often considered more efficient than the Metropolis method. However, it is important to note that in the heat bath method one must compute the transition probability for all possible changes $\Delta z^{(i)}_{k}$ in general. As a result, the computational cost scales as $O(u_{k}-l_{k})$, which can become prohibitive when the variable ranges are very wide, significantly increasing the overall runtime of SA.
@@ 3.3.3 Optimal-transition Metropolis method
Note that, if this transition were always chosen, the system would be trapped in a local minimum. To avoid this, we introduce a parameter $r=s/(N_{\text{sweep}}-1)$ with $s=0,1,\ldots,N_{\text{sweep}}-1$, defined by the progress of the annealing schedule, which controls how often the optimal transition is attempted. Namely, the probability that a candidate change $\Delta z_{k}^{(i)}$ is accepted is given by
@@
Here, $\Delta E^{\min}_{k}=\min_{\Delta z_{k}}\,\Delta E_{p,k}$ is the minimum energy difference among all possible changes of $z_{k}$. We also note that similar attempts to improve proposal mechanisms have been made in SA for continuous variables.[Dekkers1991, Ali1997, ALI200287, Locatelli2002] In general, obtaining $\Delta z^{\min}_{k}$ requires minimizing a polynomial expression of the energy difference with respect to $\Delta z_{k}$, which can be computationally expensive. However, when the polynomial degree is small, the minimum can be found efficiently. In this study, we implemented the optimal-transition Metropolis method in cases where the energy difference is at most a fourth-order polynomial in $\Delta z_{k}$, which corresponds to HUIO problems in which each variable appears with degree at most four within each monomial.
It should be noted that, unlike the Metropolis and heat bath methods, the optimal-transition Metropolis method does not satisfy the balance condition for the Boltzmann distribution even when the temperature $T$ is kept fixed. Therefore, it cannot be used for equilibrium sampling at a fixed temperature and can be employed only as a heuristic for searching low-energy states within SA.
@@ 3.4 Initial and final temperature
In this subsection, we explain how the initial and final temperatures in SA are determined.[10.1063/1.34823, Aarts1985StatisticalC] The initial temperature must be sufficiently high for the system. If the initial temperature is set too low, the system easily becomes trapped in local minima. If it is set too high, the high-temperature part of the search becomes essentially wasteful, and when the total annealing time is fixed, the low-temperature phase is effectively shortened, which again degrades the solution quality. Thus, the initial temperature should be chosen to be sufficiently high but not excessive.
The final temperature must likewise be sufficiently low for the system. If it remains too high, the energy cannot decrease enough, and the quality of the final solution is also degraded. If it is too low, on the other hand, state updates rarely occur and computation time is wasted without further improvement, although setting it too low is still preferable to setting it too high.
An important point is that the appropriate initial and final temperatures cannot be determined solely from the properties of the model. They also depend on other factors such as the total annealing time, the update scheme of the states, and the specific temperature schedule. However, it is quite difficult to determine these temperatures in a way that also accounts for these factors. In this study, therefore, we determine both temperatures solely based on the model information. To ensure robustness against variations in the annealing time or cooling schedule, the initial temperature is slightly overestimated on the high-temperature side, and the final temperature is slightly underestimated on the low-temperature side.
@@
$T_{\text{init}}=-\frac{\Delta E_{\text{max}}}{\ln a}.$ (26)
@@
$\tilde{a}^{(m)}_{k}(\bm{z})$ $=\|J_{\underbrace{k,\dots,k}_{m\text{ times}}}\|+\sum_{\begin{subarray}{c}i_{1}\\ i_{1}\neq k\end{subarray}}\|J_{i_{1},\underbrace{k,\dots,k}_{m\text{ times}}}\|z^{\text{max}}_{i_{1}}$
$+\sum_{\begin{subarray}{c}i_{1},i_{2}\\ i_{1}\neq k,i_{2}\neq k\end{subarray}}\|J_{i_{1},i_{2},\underbrace{k,\dots,k}_{m\text{ times}}}\|z^{\text{max}}_{i_{1}}z^{\text{max}}_{i_{2}}+\cdots,$ (28)
@@
$T_{\text{final}}=-\frac{\Delta E_{\min}}{\ln b}.$ (30)
@@
These definitions of the initial and final temperatures provide a practical way to set appropriate temperature ranges for SA based only on the properties of the objective function, and we use them in the same manner for all transition probabilities considered in this study.